A salesperson's weekly paycheck is 25% more than her coworker's paycheck. The two paychecks total $695. Find the amount of each paycheck.
Answer by JimboP1977(311)

Let A be the higher of the two paychecks and B the other.
A = 1.25 * B
A + B = 695
Substitute 1.25*B in for A
1.25*B + B = 695
B(1.25+1) = 695
B = 695/2.25
B = 308.89 (2 dp)
A = 1.25 * 308.89
A = 386.11 (2 dp)